# Observing dynamics of distinct structural transitions in trapped-ion clusters

### v2

@@ III.2 Buckling critical point
Figure SF3: All the out-of-plane vibrational modes of the 5-ion planar pentagon.
@@
This is a real-time video of the 6-ion cluster switching between pyramidal and bi-pyramidal configurations. The plot in Fig. 3 is obtained from this video.
Figure SF4: (Top) All the 4-ion minimum energy configurations. (Bottom) Corresponding normal-mode frequencies computed at the theoretical limit $T=0$. The grey region is the ROI for the study of mode softening at a symmetry-breaking transition.
Figure SF5: (Top) All the 5-ion minimum energy configurations. (Bottom) Corresponding normal-mode frequencies computed at the theoretical limit $T=0$. The grey region is the ROI for the study of hysteresis and triple point.

### v3

@@ II.2 Buckling critical point
Figure SF3: Toy-problem of 2 ions trapped in 1D harmonic potential. The two normal modes are c.o.m. oscillation (blue arrows) and breathing mode (green arrows). Modulation of the trap curvature leads to the driven motion along the red arrows. The Fourier spectrum of the motion shows on-resonant excitation of the breathing mode via curvature modulation.
III Collective mode excitation by modulating spring constant
Consider a toy problem of two like charges with mass m, co-trapped in a 1D harmonic potential of frequency $\omega_{t}=\sqrt{2k/m}=2\pi\times 1$ MHz as shown in Fig. SF3. The two normal modes are: c.o.m. mode with frequency $\omega_{c.o.m.}=\omega_{t}$ and breathing mode with frequency $\omega_{b}=\sqrt{3}\times\omega_{t}$. The motion of the charges is first damped to equilibrium positions, and the spring constant $k$ is then modulated at $\Omega_{m}$, i.e., $k\rightarrow k(1+\epsilon\cos(\Omega_{m}t))$. In Fig. SF3, the power spectrum of the ions’ motion after the damping period is compared with the power spectra after the modulation at different frequencies. For a modulation strength $\epsilon=10^{-3}$, the c.o.m. mode shows parametric excitation at $\Omega_{m}=2\times\omega_{t}$, while the breathing mode shows resonant excitation for $\Omega_{m}=\sqrt{3}\times\omega_{t}$. In this case, it is trivial to see that the unit vector along the driven motion overlaps is along the breathing mode (Fig. SF3).
@@ III Collective mode excitation by modulating spring constant
A video of the 6-ion cluster switching between pyramidal and bi-pyramidal configurations. The plot in Fig. 4 is obtained from this video.
Figure SF4: (Top) All the 4-ion minimum energy configurations. (Bottom) Corresponding normal-mode frequencies computed at the theoretical limit $T=0$. The grey region is the ROI for the study of Higgs-like mode softening at a symmetry-breaking transition.
Figure SF5: (Top) All the 5-ion minimum energy configurations. (Bottom) Corresponding normal-mode frequencies computed at the theoretical limit $T=0$. The grey region is the ROI for the study of hysteresis and a triple point.